OpenStudy (senia01):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y=___x+___

OpenStudy (anonymous):

Point-slope form is y-y1=m(x-x1). Plug in the given values. y-2=3(x-6) Distribute. y-2=3x-18 Add 2 to both sides. y=3x-16

OpenStudy (senia01):

Thats what i got but i was confused because when i enter the problem its a (+) and not (-)

OpenStudy (anonymous):

I see. Remember that y=3x-16 can also be written as y=3x+(-16).